QUESTION 17 A) unemployment. The type of The short term unemployment that occurs when someone is between jobs is known as the economy like technological change or when a U.S. company moves its operations to a foreign country is unemployment. Because these types of unemployment are inevitable in a large, dynamic, technologically rate of unemployment known as B) progressive, and diverse economy such as ours, the sum of these two rates is known as the C) Although it is hard to estimate, the general consensus is that it's probably somewhere bet % and E) o (enter percentage as a whole number, not a decimal). The remainder of the actual unemployment rate (most of which is due to an unemployment -referring, of course to fluctuations in the business cycle. economic contraction/recession) is called F)
avoid taxes or produce illegal goods and services in the A) economy QUESTION 16 Imagine a hypothetical economy with a population of 100 people, 80 of which over sixteen. Forty eight of these people who are working and twelve people who are willing, able and looking for work cannot find jobs. The unemployment rate in this economy is A) percentage as a whole number, not a decimal, Ho percentage signy. Suppose that 10 of those unemployed people get discouraged and give up looking for work Now, the unemployment rate is B) L (enter % (enter percentage as a whole number, not a decimal, no percentage sign QUESTION 17
QUESTION 15 The GDP may actually understate the actual l avoid taxes or produce illegal goods and services in the A) level of output in the economy by from five to thirty percent. That's because people have an incentive to economy
QUESTION 14 Imagine an economy that produces only one good, say widgets. In year 1, it produced 10 widgets widget (which sold for $10 each) and in year two in produced 8 widgets (which sold for $15 each). From year one to year 2, that economy's nominal GDP went A) enter number from sheet, not the word) and it's real GDP went B) (up/down- (up/down - enter number from sheet, not the word).
QUESTION 13 approach Because every cent that is spent will be received by somebody as income, we can compute the GDP by using the A)L or the B) approach.
Answer #1

Question 17

A) Frictional unemployment. Frictional unemployment is the time taken to find new jobs when people change jobs. There are enough job openings but it takes time to find them. This is a naturally occurring phenomena.

B) Structural unemployment. When unemployment occurs due to jobs shifting overseas, or there is a mismatch between skills required and what skills workers posses due to technological improvements.

C) Natural rate of unemployment. The natural rate of unemployment is the rate of unemployment that would occur in an economy if there were no business cyclical unemployment. It is total of frictional unemployment and structural unemployment.

D) 4.5%

E) 5%

F) Cyclical. Cyclical unemployment is unemployment due to business cycles like contraction and expansion.

Question 18

A)

Unemployment rate =(Number of unemployed/labor Force)*100

Labor force= 100-2-=80.

Unemployed=12

Employed=48

Labor force=Employed+ Unemployed=48 +12=60.

Unemployment rate= (12/60) x 100=20%.

B) New unemployed is 12-10=2

New labor force= 48+2=50

New unemployment rate= (2/50) x 100=40%.